OpenStudy (anonymous):

Fundamental Theorem of Calculus integral b=square root ofpie,a=0 5xcos(x^2) dx

OpenStudy (anonymous):

I found that F(x)=(5/2)(sin(x^2)

OpenStudy (turingtest):

and you are correct, but I don't see you using the fundamental theorem of calculus here

OpenStudy (anonymous):

So i have to find f(b)-f(a) but I know im messingup somewhere

OpenStudy (turingtest):

is the problem\[\frac d{dx}\int5x\cos(x^2)dx\]or something similar?